By substituting the value of x in the following, compute y and z

i (x)10 + (35) 8 = ( y ) 2 = ( z ) 16

ii (x)10 × (1010110) 2 = ( y ) 4 = ( z ) 10 

value of x is 49


1
Expert's answer
2021-06-30T06:01:41-0400

Part i

"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 280_{10}
\end{array}}" "\underline{\begin{array}{cc}
 X \\
 + Y
\end{array}}"

Z


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 720_{10}
\end{array}}"

121010

Hence discarding the leading 1 the result is 210.

Part ii

(x)10 × (1010110) 2 = ( y ) 4 = ( z ) 10  


(1010110)2=8610


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 086_{10}
\end{array}}" "\underline{\begin{array}{cc}
 X \\
 + Y
\end{array}}"

Z


"\underline{\begin{array}{cc}
 \space \space \space \space 490_{10} \\
 + 086_{10}
\end{array}}"

57610

Hence the result is 576